Please this is the second time I am posting, I need to learn how to set up this problem by doing dimensional analysis ONLY.

PLEASE PLEASE if you aren't going to set it up by using DIMENSIONAL ANALYSIS let someone else answer the problem. I need the steps.

  1. The healthcare provider prescribes cefazolin (Ancef) 125 mg Q6H for a child weighing 50 lb. Safe dose is 25-50 mg/kg in 4 divided doses. Available is 125 mg/mL.

Is this dose safe? Please circle one as your answer: safe, but ineffective, safe, not safe.

If safe, how many mL/dose would the nurse give? ___________mL
Answer #1

Prescibed cefazolin =125 mg QGH childs weight = 50 lb = 50/212= 22.72 Kg I recommended safe dore = 25-50 mg / kg in se divid